OpenStudy (ragingsquirrel):

What is the output value for the following function, f(x) = 5x - 2 if the input value is 3?

OpenStudy (anonymous1):

Plug in x = 3.

OpenStudy (anonymous):

13

OpenStudy (anonymous):

13

OpenStudy (anonymous1):

f(3) = 5*3 - 2 = 15 - 2 = 13.

OpenStudy (ragingsquirrel):

Thanks guys. :)